Please help me solve this equation:
solve into partial fractions
(1/(x^3+1))

Answer by stanbon(75887)

solve into partial fractions
(1/(x^3+1))
------
= 1/[(x+1)(x^2-x+1)]
-------
= A/(x+1) + (Bx+C)/(x^2-x+1)
-----
Ax^2 -Ax + A + Bx^2+Bx + Cx + C = 1
-----
Equations:
A + B = 0
-A+B+C = 0
A + C = 1
------
Solve this system of equations to get:
A = 1/3
B = -1/3
C = 2/3
=================
1/(x^3+1) = 1/(3(x+1)) + (-x+2)/[3(x^2-x+1)]
====================
Cheers,
Stan H.